ROVIBRATIONAL SPECTROSCOPY AND MOLECULAR DYNAMICS OF Ar-HBr

J. HAN, A. L. McINTOSH, Z. WANG, R. R. LUCCESE, AND \underline{J. W. BEVAN}, Department of Chemistry, Texas A and M University, College Station, TX 77843.

High resolution rovibrational spectra of several bands associated with the ν_{1} (HBr stretch) of both Ar-H^{79}Br and Ar-H^{81}Br isotopomers have been recorded using a tunable infrared diode laser supersonic jet spectrometer. In addition to the ν_{1} fundamental, the combination bands ν_{1} + 2ν_{2}^{0}, ν_{1} + ν_{2}^{1}, ν_{1} + ν_{3}, ν_{1} + 2ν_{2}^{0} + ν_{3}, the hot band ν_{1} +2ν_{2}^{0} −2ν_{2}^{0}, and the difference band ν_{1} −2ν_{2}^{0} were observed and analyzed. Nuclear quadrupole splittings have been partially resolved in Q branch transitions of the ν_{1} + ν_{2}^{1} band. A strong Coriolis perturbation between the ν_{1} +ν_{2}^{1e} and ν_{1} +ν_{3} was observed and analyzed. These results have been incorporated to determine an improved H4 intermolecular potential for Ar-HBr. In addition, the intermolecular potential surface for the Ar-HBr with HBr in its v = 1 state has also been modeled based on the experimental results of this work.